The total earnings of a worker are represented by E = 100 + $10(24−L), where E is earnings and L is the number of hours of leisure. How many hours of leisure are consumed if this worker's total earnings are $160? Show solutions.


1
Expert's answer
2022-02-10T13:48:46-0500

SOLUTION

The equation of total earnings of a worker is E = 100 + $10(24−L)


But we have been given that the earnings of a worker are $160


We substitute the worker earnings in our equation as shown below;


160 = 100 + $10(24−L)

160 = 100 + 240 - 10L

160 =340 -10L

160 - 340 = -10L

-180 = -10L

18 = L


ANSWER

The worker consumes 18 hours of leisure